# Nudging Nutrition: Lessons from the Danish “Fat Tax”

### v2

@@ 3.2 Identification strategy, estimation, and inference
Some assumptions that are standard in literature need to be satisfied for Equations (1) to (3) to be valid. The first assumption rules out anticipation effects. This implies that Danish households did not adjust their consumption behavior in anticipation of the tax, e.g. by hoarding. As the tax was adopted approximately half a year before its implementation, we cannot exclude the presence of hoarding. We rigorously study the presence of anticipation effects of the fat tax in Section 4.1.
The second assumption is the parallel trends assumption. This assumption states that in the absence of the Danish fat tax, the outcome trends in Denmark would have followed the same trends as the ones observed in Northern-Germany during the tax period. If the composition of the treatment and control groups are different in terms of household characteristics, then this might cause the German and Danish trends to diverge in case of heterogeneous treatment effects, leading to a violation of the unconditional parallel trends assumption (see Roth et al., (2023)). To address this concern, we relax the parallel trends assumption slightly and allow for the presence of covariate-specific trends by conditioning on pre-tax household characteristics. When controlling for observed covariates we also need to impose the standard overlap assumption, i.e., it must hold that for all $X$ that $p(X)<0$. This assumption ensures that for each Danish household, we can find German households with similar characteristics. We investigate the plausibility of the parallel trends and overlap assumptions in Section 4.1.
@@
In Section 5 we run further robustness checks. First, we exclude Danish households living close to the German border, since the introduction of a tax can lead to increased border trade to avoid the tax. Second, we investigate how much our results rely on controlling for household characteristics by estimating the effects unconditionally.
@@ 4 Results
In this section, we report the main results. [^11] We tailored the R package of Callaway and Sant’Anna, (2021) to our setting and express our gratitude to the original authors for making their code publicly available.
@@ 4.1 Plausibility of identifying assumptions
We first examine the plausibility of the no anticipation assumption which is crucial for DID. As mentioned in Section 3.2, this assumption posits that households did not change their behavior in anticipation of the tax implementation. To explore this, we estimate the effect of the tax in the two quarters preceding its introduction, namely the second and third quarters of 2011. The tax was announced on March 17, 2011, and introduced in October, 2011. The results reported in Table 1 do not show any statistically significant anticipation effect. Nevertheless, we take a conservative stand and, to avoid any potential bias due to anticipation effects, we exclude the second and third quarter of 2011 from the analysis. We find qualitatively similar results including these two quarters. [^12] The results are available from the authors upon request.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml $ATT_{2011-Q2}$13.910 82.603 -142.069 -45.696 38.916 -54.424 54.246 106.351 (se) (47.829) (107.047) (160.676) (94.261) (60.690) (124.059) (47.474) (51.176) $ATT_{2011-Q3}$ 26.963 95.758 -99.659 -37.812 45.896 202.126 76.975 72.012 (se) (42.933) (100.254) (143.327) (87.195) (63.777) (135.263) (49.936) (48.075) Panel B: Quarterly average expenditure in Euro cents $ATT_{2011-Q2}$19.721 14.544 -48.137 -13.068 32.325 -44.864 71.076 28.082 (se) (39.060) (49.050) (130.123) (28.545) (38.540) (28.308) (48.302) (14.516) $ATT_{2011-Q3}$ 8.636 51.837 -12.612 -4.428 12.671 39.789 99.147 20.986 (se) (36.968) (47.847) (133.483) (28.970) (40.789) (30.300) (47.703) (13.646) Panel C: Quarterly average amount (in packages) consumed $ATT_{2011-Q2}$ 0.009 0.243 -0.145 -0.309 0.087 -0.095 0.317 0.263 (se) (0.231) (0.419) (0.497) (0.366) (0.255) (0.269) (0.296) (0.207) $ATT_{2011-Q3}$ 0.182 0.196 -0.278 -0.200 0.059 0.405 0.522 0.197 (se) (0.188) (0.427) (0.484) (0.320) (0.270) (0.283) (0.279) (0.185) Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors are clustered at household level. The results are obtained controlling for the age of the head of the household, the number of children under 15 years, the gender of diary keeper, the educational level (ISCED), the income level, and the household size.
A second aforementioned crucial assumption for the credibility of our estimates is the parallel trends assumption. An advantage of having multiple pre-tax time periods is the opportunity to gain some insight into the plausibility of this assumption. We assess the plausibility in two different ways. First, Table 2 contains the p-values of a Wald test for parallel trends, where the assumption of parallel trends in the pre-tax period is tested under the null hypothesis that they are parallel. Notably, the p-values derived from this approach do not raise concerns, as we fail to reject the null hypothesis of parallel trends, with the exception of butter in the expenditure outcome. Second, the results of pseudo treatment effects for the pre-tax periods are shown in Appendix C. Reassuringly, the uniform confidence intervals in all pre-tax periods contain zero, offering further evidence in support of the parallel trends assumption for all product categories and all outcomes. Altogether, inspection of the pre-tax trends lends strong support of the plausibility of parallel trends between Danish and Northern-German households in the tax period.
@@ 4.2 The causal effect of the fat tax on consumer responses during the tax period
The results depicted in Panel A of Table 2 reveal a statistically significant decrease in the consumption of cheese, liver sausage, and salami during the tax period. Specifically, there is an average reduction of approximately 13% for cheese, 11% for liver sausage, and 12% for salami compared to the pre-tax quarter. While we do not observe significant changes in expenditure for cheese and salami, there was a noteworthy 13.75% (about 0.8 Euro) reduction in expenditure for liver sausage, as illustrated in Table 2, Panel B. Conversely, we observe a substantial increase in expenditure for butter (2.34 Euro, 25.2%), cream (0.7 Euro, 14.49%), margarine (0.8 Euro, 15.42%), and sour cream (0.27 Euro, 16%) despite the absence of any discernible impact on the consumption of these products.
@@
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) $-56.891$ 73.292 $-351.109^{*}$ 27.663 $-120.925^{*}$ 114.433 $-66.774^{**}$ 27.768 (se) (35.788) (78.927) (127.263) (65.761) (47.718) (101.072) (36.823) (36.351) % change -10.07% 4.81% -13.05% 2.01% -11.07% 6.24% -12.10% 4.85% obs. 29,645 40,929 32,129 39,274 31,246 38,687 31,243 36,121 p-val. pre-trend 0.273 0.384 0.344 0.455 0.846 0.772 0.963 0.208 Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) -25.075 $234.879^{***}$ $-240.166$ $70.823^{**}$ $-81.437^{*}$ $84.643^{***}$ $-50.452$ $27.131^{*}$ (se) (26.216) (37.992) (105.520) (19.589) (28.147) (22.041) (38.160) (10.413) % change -6.85% 25.20% -10.42% 14.49% -13.75% 15.42% -8.81% 16.03% obs. 29,645 40,929 32,129 39,274 31,246 38,687 31,243 36,121 p-val. pre-trend 0.653 0.000 0.505 0.748 0.929 0.603 0.921 0.075 Panel C: Quarterly average amount (in packages) consumed Danish fat tax (ATT) -0.010 0.361 -0.304 -0.202 $-0.463$ $0.404$ -0.009 0.059 (se) (0.160) (0.314) (0.398) (0.272) (0.199) (0.198) (0.233) (0.145) % change -0.51% 6.01% -7.09% -6.19% -14.40% 10.39% -0.33% 4.77% obs. 29,645 40,929 32,129 39,274 31,246 38,687 31,243 36,121 p-val. pre-trend 0.987 0.557 0.249 0.954 0.984 0.795 0.977 0.512 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
Notably, there is no observed statistically significant impact on the quantity of packages purchased across all products. Nevertheless, the observed decrease in consumption of cheese, liver sausage, and salami hints at a potential adjustment by producers who may have responded to the imposed tax by reducing the package size of these three items. In contrast, our results suggest that the package sizes for butter, cream, margarine, and sour cream remained unaltered, as there was no significant impact on either the number of packages purchased or the weight consumed. At the same time, we observe a significant rise in expenditure, possibly driven by elevated prices and a low elasticity of demand for these particular products.
In summary, our findings indicate that the tax successfully achieved the intended reduction in the consumption of cheese, liver sausage, and salami during the tax period. However, it did not yield the desired effect on the consumption of the other products under examination. Notably, the tax led to increased consumer expenditure on butter, cream, margarine, and sour cream without significantly affecting the consumption of these particular products.
@@ 4.3 The causal effect of the fat tax on consumer responses after its abolishment
Table 3: The causal effect of the Danish fat tax on the average quarterly consumption and expenditure during the two years after abolishment of the tax (January 2013 to December 2014), using the final tax period (October 2012 to December 2012) as reference period.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) $124.442^{**}$ 5.298 $557.856^{***}$35.433 $195.790^{***}$ 72.082 $131.118^{**}$ 78.072 (se) (42.177) (86.445) (115.098) (89.103) (50.680) (98.307) (40.289) (40.092) % change 23.43% 0.33% $2$54% 19.36% 4.20% 25.39% 13.17% obs. 22,354 28,507 23,735 27,544 23,315 27,085 23,312 25,831 p-val. pre-trend 0.310 0.000 0.296 0.335 0.334 0.018 0.414 0.564 Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) 57.635 $-171.538^{***}$ $415.404^{***}$24.235 $122.264^{**}$ $-67.918^{**}$ $136.449^{***}$ 7.812 (se) (29.636) (42.352) (92.873) (26.516) (33.875) (22.868) (38.423) (11.110) % change 15.85% -16.77% 19.19% -4.51% 22.67% -13.20% 25.21% 4.28% obs. 22,354 28,507 23,735 27,544 23,315 27,085 23,312 25,831 p-val. pre-trend 0.456 0.000 0.358 0.833 0.051 0.013 0.252 0.093 Panel C: Quarterly average amount (in packages) consumed Danish fat tax (ATT) $0.430^{**}$0.017 $0.938^{*}$ 0.145 $0.782^{**}$ 0.144 $0.654^{**}$ 0.341 (se) (0.158) (0.340) (0.393) (0.327) (0.255) (0.187) (0.234) (0.175) % change 21.53% -0.26% 21.14% 4.47% 26.19% 3.87% 23.69% 26.58% obs. 22,354 28,507 23,735 27,544 23,315 27,085 23,312 25,831 p-val. pre-trend 0.128 0.000 0.914 0.956 0.577 0.002 0.881 0.495 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
@@
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) 0.586 $256.953^{**}$ 31.182 117.048 1.505 $325.944^{**}$ 5.839 $90.648^{*}$ (se) (35.004) (72.408) (129.746) (71.148) (46.676) (92.464) (37.856) (35.617) % change 0.10% 16.05% 1.15% 8.31% 0.14% 18.05% 1.05% 15.41% obs. 33,510 45,295 36,043 43,648 35,160 42,954 35,175 40,299 p-val. pre-trend 0.182 0.362 0.268 0.369 0.859 0.853 0.959 0.138 Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) 0.024 $162.860^{***}$ 12.967 $57.764^{**}$33.561 39.009 21.002 $37.947^{**}$ (se) (26.678) (38.977) (105.804) (19.152) (30.410) (21.705) (38.239) (10.451) % change 0.01% 16.92% 0.56% 11.39% -5.73% 7.66% 3.59% 21.65% obs. 33,510 45,295 36,043 43,648 35,160 42,954 35,175 40,299 p-val. pre-trend 0.516 0.000 0.235 0.668 0.950 0.734 0.905 0.050 Panel C: Quarterly average amount (in packages) consumed Danish fat tax (ATT) 0.175 $1.348^{***}$ 0.586 0.058 -0.001 $0.916^{***}$ $0.597^{*}$ 0.285 (se) (0.150) (0.294) (0.368) (0.254) (0.196) (0.210) (0.222) (0.145) % change 8.53% 20.86% 12.83% 1.74% -0.04% 23.65% 21.17% 22.28% obs. 33,510 45,295 36,043 43,648 35,160 42,954 35,175 40,299 p-val. pre-trend 0.982 0.472 0.282 0.939 0.991 0.797 0.989 0.481 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
The fat tax was abolished in January 2013. We study the two years after the abolishment to gain insight into potential habit formation induced by the tax. We present the effects of the abolishment of the fat tax on consumption in comparison to the final quarter of the tax period in Table 3 and in comparison to the pre-tax period in Table 4.
For cheese, liver sausage, and salami, we find that consumption increased after the tax was abolished, as shown in Panel A of Table 3. This increase brought the consumption of these products back to pre-tax levels, as indicated in Panel A of Table 4. Regarding salami, we observe that the abolishment of the tax resulted in a consistent reduction in packaging size, with a 23.69% increase in the average number of packages purchased. However, this change had no significant effect on the average consumption in terms of weight. Overall, despite the short-term reduction in consumption observed during the tax period, these effects seemed to have disappeared after the tax was abolished. This is in line with Allais et al., (2010), who found that long-term effects can take seven years to establish, as mentioned in the literature review.
Panel A of Table 4 shows that butter, margarine and sour cream experienced a substantial increase in consumption after the abolishment of the tax in comparison to the pre-tax period as well as an overall increase in expenditure and average number of packages purchased (see Panel B and C).
In summary, the tax failed in creating any long-term reductions in consumption of the targeted products, and its abolishment induced a short-term increase in consumption for some of these products.
4.4 Price mechanism

### v3

@@ 3.3 Causal Effects Definition and Estimation
We adopt the estimator for event-study designs proposed by Callaway and Sant’Anna, (2021), which extends the doubly robust estimator of Sant’Anna and Zhao, (2020) to the setting with multiple time periods. [^12] Note that Callaway and Sant’Anna, (2021) additionally consider the setting with variation in treatment timing (staggered adoption), where individuals belong to a certain group depending on the timing of their treatment. However, in our context, the number of groups equals one as the tax was simultaneously introduced and repealed for all individuals.  [^13] We tailored the R package of Callaway and Sant’Anna, (2021) to our setting and express our gratitude to the original authors for making their code publicly available. The main advantage of this approach over other estimators, such as the common two-way fixed effects estimator (see, for instance, Roth et al., (2023) for a recent review of the DID literature), is that it allows to control for observed covariates in a flexible way.
@@
Some assumptions that are standard in literature need to be satisfied for Equations (1) to (3) to be valid. The first assumption rules out anticipation effects. This implies that Danish households did not adjust their consumption behavior in anticipation of the tax, e.g., by hoarding. As the tax was adopted approximately half a year before its implementation, we cannot exclude the presence of hoarding. We rigorously study the presence of anticipation effects of the fat tax in Section 4.1.
The second assumption is the parallel trends assumption. This assumption states that in the absence of the Danish fat tax, the outcome trends in Denmark would have followed the same trends as the ones observed in Northern-Germany during the tax period. If the composition of the treatment and control groups are different in terms of household characteristics, then this might cause the German and Danish trends to diverge in case of heterogeneous treatment effects, leading to a violation of the unconditional parallel trends assumption (Roth et al., , 2023). To address this concern, we relax the parallel trends assumption slightly and allow for the presence of covariate-specific trends by conditioning on pre-tax household characteristics. When controlling for observed covariates we also need to impose the standard overlap assumption, i.e., it must hold that for all $X$ that $p(X)<0$. This assumption ensures that for each Danish household, we can find German households with similar characteristics. We investigate the plausibility of the parallel trends and overlap assumptions in Section 4.1.
@@
In Section 5, we describe the robustness checks we have performed. First, we exclude Danish households living close to the German border, since the introduction of a tax can lead to increased border trade to avoid the tax. Second, we investigate how much our results rely on controlling for household characteristics by estimating the effects unconditionally.
@@ 4.1 Plausibility of identifying assumptions
We first examine the plausibility of the no anticipation assumption which is crucial for DID. As mentioned in Section 3.1, this assumption posits that households did not change their behavior in anticipation of the tax implementation. To explore this, we estimate the effect of the tax in the two quarters preceding its introduction, namely the second and third quarters of 2011. The tax was announced on March 17, 2011, and introduced in October, 2011. The results reported in Table 10 in Appendix C do not show any statistically significant anticipation effect. Nevertheless, we take a conservative stand and, to avoid any potential bias due to anticipation effects, we exclude the second and third quarter of 2011 from the analysis. We find qualitatively similar results including these two quarters. [^14] The results are available from the authors upon request.
A second aforementioned crucial assumption for the credibility of our estimates is the parallel trends assumption. An advantage of having multiple pre-tax time periods is the opportunity to gain some insight into the plausibility of this assumption. We assess the plausibility in two different ways. First, Table 1 contains the p-values of a Wald test for parallel trends, where the assumption of parallel trends in the pre-tax period is tested under the null hypothesis that they are parallel. Notably, the p-values derived from this approach do not raise concerns, as we fail to reject the null hypothesis of parallel trends, with the exception of butter in the expenditure outcome. Second, the results of pseudo treatment effects for the pre-tax periods are shown in Appendix D. Reassuringly, the uniform confidence intervals in all pre-tax periods contain zero, offering further evidence in support of the parallel trends assumption for all product categories and all outcomes. Altogether, inspection of the pre-tax trends lends strong support of the plausibility of parallel trends between Danish and Northern-German households in the tax period.
@@
4.2 Two-way fixed effects estimates
The Two-Way Fixed Effects (TWFE) estimates reported in Table 11 in Appendix C indicate that the Danish fat tax had differential impacts across various high-fat food products. The tax effectively reduced consumption and expenditure for items like bacon, cheese, liver sausage, and salami, demonstrating a significant decrease in their intake. Conversely, for products such as butter and sour cream, expenditure increased despite minor reductions in consumption, suggesting inelastic demand where consumers were willing to pay higher prices without substantially altering their purchasing habits.
@@ 4.3 The causal effect of the fat tax on consumer responses during the tax period
As mentioned in the introduction, we employ the Callaway and Sant’Anna, estimator in our main specification instead of the traditional TWFE model. This choice is motivated by the estimator’s ability to provide more robust inference and to flexibly incorporate covariates, thereby enhancing the reliability and validity of our results, particularly in the presence of treatment effect heterogeneity.
The results depicted in Panel A of Table 1 are generally consistent with those obtained from the TWFE estimates but with some notable differences. Specifically, we observe a statistically significant decrease in the consumption of cheese and liver sausage during the tax period. There is an average reduction of approximately 13.36% for cheese (p-value = 0.058) and 10.56% for liver sausage (p-value = 0.019) compared to the pre-tax quarter. The decrease in cheese consumption is significant at the 10% level, while the reduction in liver sausage consumption is significant at the 5% level.
For salami, although there is a reduction of 14.38%, the p-value is 0.270, indicating that this decrease is not statistically significant. Similarly, we do not observe significant changes in the consumption of bacon, butter, cream, margarine, or sour cream during the tax period.
In Panel B of Table 1, we see substantial increases in expenditure for butter, cream, margarine, and sour cream, despite the absence of any discernible impact on their consumption. Specifically, expenditures rose by 2.35 euro cents (25.20%) for butter (p-value = 0.000), 0.71 (14.49%) for cream (p-value = 0.019), 0.85 euro cents (15.42%) for margarine (p-value = 0.005), and 0.27 euro cents (16.03%) for sour cream (p-value = 0.092). These increases are statistically significant, with butter and margarine significant at the 1% level, cream at the 5% level, and sour cream at the 10% level.
Conversely, we observe a decrease in expenditure for liver sausage by 0.55 euro cents (9.19%), which is marginally significant (p-value = 0.052). This aligns with the decreased consumption of liver sausage. Expenditure decreases for bacon and salami are not statistically significant, indicating that the tax did not significantly affect spending on these products.
These results suggest that the tax led to higher prices for butter, cream, margarine, and sour cream, which consumers were willing to pay without significantly reducing their consumption—indicating inelastic demand for these products. In contrast, the tax effectively reduced both the consumption and expenditure of cheese and liver sausage.
@@
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) $-64.363$ 73.292 $-345.752^{*}$ 27.663 $-115.680^{**}$ 114.433 $-54.267$ 27.768 (se) (31.976) (81.409) (127.169) (63.224) (33.229) (95.311) (34.480) (37.123) p-value 0.166 0.536 0.058 0.755 0.019 0.414 0.270 0.602 % change $-11.39$% 4.81% $-13.36$% 2.01% $-10.56$% 6.24% $-14.38$% 4.85% obs. 28,874 40,929 32,020 39,274 29,417 38,687 30,800 36,121 p-val. pre-trend 0.092 0.384 0.333 0.455 0.196 0.772 0.978 0.208 Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) $-26.947$ $234.879^{***}$ $-243.321^{*}$ $70.823^{**}$ $-54.601^{*}$ $84.643^{***}$ $-53.808$ $27.131^{*}$ (se) (22.878) (36.546) (104.101) (21.156) (19.624) (20.909) (34.438) (10.745) p-value 0.430 0.000 0.086 0.019 0.052 0.005 0.275 0.092 % change -7.36% 25.20% -10.91% 14.49% -9.19% 15.42% -11.98% 16.03% obs. 28,874 40,929 32,020 39,274 29,417 38,687 30,800 36,121 p-val. pre-trend 0.175 0.000 0.501 0.748 0.304 0.603 0.994 0.075 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
Our findings highlight the nuanced effects of the fat tax across different products. While the tax was effective in reducing consumption of certain high-fat items like cheese and liver sausage, it did not achieve the same outcome for other products such as butter and cream. Instead, consumers continued purchasing these goods despite higher prices, leading to increased expenditures. This underscores the importance of considering demand elasticities and potential substitution effects when designing taxation policies aimed at altering consumer behavior to achieve public health objectives.
@@ 4.4 The causal effect of the fat tax on consumer responses after its abolishment
The fat tax was abolished in January 2013. To gain insight into potential habit formation induced by the tax, we study the two years following its abolishment, comparing consumption and expenditure to the pre-tax period. We present the effects of the abolishment of the fat tax on consumption and expenditure in Table 2.
For butter, margarine, and sour cream, we find that consumption significantly increased after the tax was abolished, as shown in Panel A of Table 2. Specifically, butter consumption increased by 256.95 grams per quarter (a 16.05% increase, p-value = 0.023), margarine consumption increased by 325.94 grams per quarter (an 18.05% increase, p-value = 0.030), and sour cream consumption increased by 90.65 grams per quarter (a 15.41% increase, p-value = 0.083).
These increases in consumption were accompanied by significant increases in expenditure for these products, as shown in Panel B of Table 2. Expenditure on butter increased by 162.86 euro cents (16.92%, p-value = 0.003), and expenditure on sour cream increased by 37.95 euro cents (21.65%, p-value = 0.010). Expenditure on cream also showed an increase of 57.76 euro cents (11.39%, p-value = 0.055), which is marginally significant.
For other products such as cheese, liver sausage, and salami, we do not observe significant changes in consumption or expenditure after the tax was abolished. The ATT estimates for these products are not statistically significant, indicating that consumption levels returned to or remained at pre-tax levels.
Overall, despite the short-term reduction in consumption observed during the tax period for some products, these effects did not persist after the tax was abolished. Instead, the consumption of certain high-fat products increased, surpassing pre-tax levels in some cases. This suggests that the tax did not lead to lasting changes in consumer behavior.
These findings are consistent with previous research, such as Allais et al., (2010), which indicates that long-term effects of food taxes on consumption can take several years to establish. In our study, the abolishment of the tax led to a rebound in consumption, highlighting the challenges in achieving sustained dietary changes through taxation alone.
In summary, the fat tax failed to create any long-term reductions in the consumption of the targeted products. Its abolishment was followed by significant increases in consumption and expenditure for products like butter, margarine, and sour cream. This underscores the importance of considering not only the immediate impact of such taxes but also their long-term effectiveness.
@@
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) -18.080 $256.953^{**}$ 29.770 117.048 -5.385 $325.944^{**}$ 19.274 $90.648^{*}$ (se) (30.325) (77.633) (128.601) (69.297) (33.609) (101.442) (33.354) (34.894) p-value 0.692 0.023 0.874 0.242 0.914 0.030 0.699 0.083 % change -3.13% 16.05% 1.14% 8.31% -0.49% 18.05% 4.90% 15.41% obs. 32,648 45,295 35,949 43,648 33,178 42,954 34,711 40,299 p-val. pre-trend 0.051 0.362 0.251 0.369 0.168 0.853 0.980 0.138 Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) 3.263 $162.860^{***}$ 9.141 $57.764^{*}$3.958 39.009 7.371 $37.947^{**}$ (se) (24.635) (36.912) (98.288) (20.108) (19.625) (22.154) (33.382) (9.741) p-value 0.928 0.003 0.949 0.055 0.892 0.229 0.880 0.010 % change 0.86% 16.92% 0.40% 11.39% -0.67% 7.66% 1.59% 21.65% obs. 32,648 45,295 35,949 43,648 33,178 42,954 34,711 40,299 p-val. pre-trend 0.124 0.000 0.230 0.668 0.212 0.734 0.995 0.050 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
4.5 Producer responses
4.5.1 Price mechanism